# R.A.C.E. : Robust Adversarial Concept Erasure for Secure Text-to-Image Diffusion Model

### v1

@@ Abstract
In the evolving landscape of text-to-image (T2I) diffusion models, the remarkable capability to generate high-quality images from textual descriptions faces challenges with the potential misuse of reproducing sensitive content. To address this critical issue, we introduce Robust Adversarial Concept Erase (RACE), a novel approach designed to mitigate these risks by enhancing the robustness of concept erasure method for T2I models. RACE utilizes a sophisticated adversarial training framework to identify and mitigate adversarial text embeddings, significantly reducing the Attack Success Rate (ASR). Impressively, RACE achieves a 30 percentage point reduction in ASR for the “nudity” concept against the leading white-box attack method. Our extensive evaluations demonstrate RACE’s effectiveness in defending against both white-box and black-box attacks, marking a significant advancement in protecting T2I diffusion models from generating inappropriate or misleading imagery. This work underlines the essential need for proactive defense measures in adapting to the rapidly advancing field of adversarial challenges.
@@ 1 Introduction
In response to the challenges posed by the malicious exploitation of generative models, Stable Diffusion (SD)[46] has integrated a safety checker[45] and advocates for the utilization of a watermarking module [71]. Despite these initiatives, the reliance on post-hoc interventions presents limitations due to their potential for circumvention [23, 8]. Consequently, the research community is pivoting towards formulating methodologies that embed safety protocols directly within the image generation pipeline. This approach ensures that content security is an intrinsic aspect of image creation [59, 66, 67, 22, 37, 23, 8].
Although these methods are valuable for identifying the source of content after an incident, their reactive nature highlights the necessity for proactive strategies. This includes pioneering techniques attempting early removal of sensitive content from T2I models [28, 10], thereby preempting the production of inappropriate or harmful material.
To address the challenge of removing undesirable content from T2I models, even when users attempt to circumvent content restrictions, recent research has focused on the development of concept erasure techniques within T2I diffusion models [10, 28, 69, 13]. These techniques primarily aim to eliminate specific concepts (e.g., “nudity”) by altering the text embeddings associated with these concepts to neutral representations. Despite these efforts, there remains a vulnerability wherein erased concepts can be reconstructed. This is achieved by identifying text tokens that closely align with the visual embeddings of the targeted concepts, thus enabling the regeneration of prohibited content [4, 72, 55, 60, 64]. This issue is illustrated in Fig. 1, which demonstrates that even with concept erasure, T2I models can be manipulated through prompt modification to regenerate the restricted content. This underscores the imperative for a more robust concept erasure methodology that can withstand such reconstruction attempts, ensuring the integrity of content generation within T2I models.
@@ 3.2 Adversarial Training on Concept Erased Diffusion Models
To assess the efficacy of our method, we execute tests on a $\Phi_{\theta}$ model trained for the removal of the “nudity” concept via the ESD method. Utilizing 142 nudity-centric prompts from the I2P dataset [51], we systematically select timesteps $t^{*}$ at 100-step intervals within the $[1,1000]$ range and compute the Attack Success Rate (ASR), as formulated in Eq.(8), for each $t^{*}$. The results, depicted in Fig. 2, unveil a notable capability: the reconstruction of the erased concept is feasible even with the attack confined to a single timestep $t^{*}$. Interestingly, after $t^{*}=500$, we observe an increasing trend in the ASR. The modified images resulting from the attack visually convey the transition back to the erased concept and reflect the ASR patterns. For additional insights into various concepts and their corresponding ASR trends across different timesteps $t^{*}$, the supplementary material offers further information.
@@
Adversarial Training for T2I Concept Erasure. Motivated by the findings from our single-timestep adversarial attack experiments, we explore the potential of such attacks to enhance the robustness of concept erasure in T2I models, posing the question: “Can adversarial attacks improve the resilience of concept erasure mechanisms?”
@@ 4.1 Experimental Setting
Table 2: Attack Success Rate (ASR) against white (○)/black (●) box attacks. We conduct experiments on artistic style (Van Gogh), explicit concepts (nudity, violence, and illegal act), and objects (church, golf ball, and parachute). “ESD/RACE-Concept” denotes the concept erased from the model. We also measure the quality of T2I models. We can observe that RACE reduces ASR by over 30% (-0.30 or below) in absolute value on Van Gogh, nudity, and church for UnlearnDiff [72], which is the previous SOTA attack method.
@@ 4.3 Disentanglement
Figure 4: RACE’s Disentanglement in Concept Erasure. This figure highlights RACE’s precision in erasing specific concepts, as shown in diagonal images, while preserving unrelated concepts, evident in off-diagonal images. For reference, baseline images generated by the original Stable Diffusion (SD) model are also presented.
@@
Investigating RACE’s disentanglement performance, our study focuses on its capability to precisely erase intended concepts without impacting other elements. The evaluation spans both qualitative and quantitative measures.
On the qualitative front, we configure separate RACE-enhanced models to specifically erase artistic imprints such as “Van Gogh”, “Thomas Kinkade”, and “Kilian Eng” from the SD. The illustrative outcomes, showcased in Fig. 4, underscore RACE’s erasure precision, ensuring that the excision of one artistic style doesn’t lead to the collateral removal of others. This meticulous erasure extends to discrete object concepts like “church”, “golf ball”, and “parachute”, with the generated images further affirming the method’s discernment.
@@ 4.4 Discussion
Table 4: Ablation Studies for Improving Performance. Using the “nudity” concept, we conduct effectiveness by adding weight regularization or close concept keywords.
@@
Potential Strategy to Improve the Robustness-Quality Tradeoff. Our findings underscore the ability of RACE to significantly bolster the SD model’s defense against prompt-based adversarial attacks across a variety of concepts. However, as illustrated in Tab.2, enhancing robustness appears to inversely impact image quality. To address this dichotomy, we test a refined version of the RACE loss function incorporating a regularization term:
@@ 5 Conclusion
In this work, we present RACE, a novel defense approach designed to protect the Text-to-Image Stable Diffusion models from prompt-based red teaming attacks. RACE effectively strengthens the model’s concept erasure capabilities while maintaining computational efficiency, offering a valuable enhancement to the current erasure framework and bolstering defenses against various adversarial techniques. We also observe the robustness-quality tradeoff and discuss possible future directions to improve it. This initial contribution lays the groundwork for further exploration, underscoring the critical importance of developing sophisticated defenses in the rapidly evolving domain of generative AI.
6 Acknowledgement
We would like to express our gratitude to Sangmin Jung for assisting with the experiments. This work is partially supported by the National Science Foundation under Grant No. 2038666 and No. 2101052. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
Appendix A Additional Related Works
A.1 Adversarial Training Approaches
@@
A.2 Additional Works in Concept Erase
@@
Appendix B Additional Training Details
@@
Appendix C Additional Analysis of Single-Timestep Adversarial Attack
In order to demonstrate the effectiveness of the proposed single-timestep adversarial attack, we perform additional experiments on $\Phi_{\theta}$ models each trained to erase the “Van Gogh” and “church” concepts via the ESD method [10]. The results, as illustrated in 5, demonstrate that a single-timestep adversarial attack can successfully reconstruct previously erased concepts.
Appendix D Extended Visual Results from RACE
In Figures 6, 7, and 8, we present a supplementary collection of images to further demonstrate the capability of our method in erasing targeted concepts. It is important to note that the set of images in 7 contains explicit content.
@@ Appendix D Extended Visual Results from RACE
Appendix E Determining PGD Hyperparameters
Within the main paper, we settle on $\epsilon=0.1$ and designate 10 steps for the adversarial attack. This section delves into the empirical analysis underpinning this selection, particularly focusing on the balance between ASR efficacy and hyperparameter tuning. As delineated in 5, an $\epsilon$ value of 0.2 marginally outperforms 0.1. However, as elaborated in the main manuscript, an elevated attack intensity with $\epsilon=0.2$ may compromise the equilibrium between adversarial robustness and the fidelity of generated images. Holding $\epsilon$ at 0.1, we observe minimal variance between 10 and 20 attack steps, reinforcing our decision to configure PGD parameters at $\epsilon=0.1$ and 10 steps for an optimal trade-off.
@@ Appendix E Determining PGD Hyperparameters
Appendix F Challenges in Erasing Violence and Illegal Act
The data presented in Table 2 of the main manuscript indicate that “violence” and “illegal act” concepts exhibit less pronounced reductions in Attack Success Rate (ASR) against UnlearnDiff white box attack [72]. Specifically, “nudity” sees a 30% decrease in ASR, while “violence” and “illegal act” experience reductions of 11% and 5%, respectively. To align with the experimental framework of UnlearnDiff [72], we employed “violence” and “illegal act” as keywords for concept removal. It is postulated that these terms may not optimally represent the range of hazardous imagery producible by Stable Diffusion (SD) [46] through the prompts in the I2P dataset [51]. To examine this hypothesis, we calculate the CLIP-score [14] for explicit images generated by a baseline SD model, $SD(\tilde{y})$, where $\tilde{y}$ refers prompts from I2P and $c$ includes “nudity”, “violence”, and “illegal act” (i.e. CLIP-score($SD(\tilde{y})$,$c$)). As illustrated in 6, a higher CLIP-score, indicating better keyword alignment with the dangerous images, correlates with a more substantial decrease in ASR. Conversely, keywords that poorly match the hazardous image content tend to result in less significant ASR reductions. This underscores the critical importance of selecting highly representative keywords for concept erasure in T2I models, as their alignment with the generated content significantly influences the effectiveness of concept erase.

### v2

@@ Abstract
In the evolving landscape of text-to-image (T2I) diffusion models, the remarkable capability to generate high-quality images from textual descriptions faces challenges with the potential misuse of reproducing sensitive content. To address this critical issue, we introduce Robust Adversarial Concept Erase (RACE), a novel approach designed to mitigate these risks by enhancing the robustness of concept erasure method for T2I models. RACE utilizes a sophisticated adversarial training framework to identify and mitigate adversarial text embeddings, significantly reducing the Attack Success Rate (ASR). Impressively, RACE achieves a 30 percentage point reduction in ASR for the “nudity” concept against the leading white-box attack method. Our extensive evaluations demonstrate RACE’s effectiveness in defending against both white-box and black-box attacks, marking a significant advancement in protecting T2I diffusion models from generating inappropriate or misleading imagery. This work underlines the essential need for proactive defense measures in adapting to the rapidly advancing field of adversarial challenges. Our code is publicly available: https://github.com/chkimmmmm/R.A.C.E.
Keywords:
Concept Erasure Responsible Image Generative Models Secure T2I Diffusion Models
@@ 1 Introduction
In response to the challenges posed by the malicious exploitation of generative models, Stable Diffusion (SD)[46] has integrated a safety checker[45] and advocates for the utilization of a watermarking module [71]. Despite these initiatives, the reliance on post-hoc interventions presents limitations due to their potential for circumvention [22, 8]. Consequently, the research community is pivoting towards formulating methodologies that embed safety protocols directly within the image generation pipeline. This approach ensures that content security is an intrinsic aspect of image creation [60, 67, 66, 23, 37, 22, 8]. Although these methods are valuable for identifying the source of content after an incident, their reactive nature highlights the necessity for proactive strategies. This includes pioneering techniques attempting early removal of sensitive content from T2I models [27, 10], thereby preempting the production of inappropriate or harmful material.
To address the challenge of removing undesirable content from T2I models, even when users attempt to circumvent content restrictions, recent research has focused on the development of concept erasure techniques within T2I diffusion models [10, 27, 69, 13]. These techniques primarily aim to eliminate specific concepts (e.g., “nudity”) by altering the text embeddings associated with these concepts to neutral representations. Despite these efforts, there remains a vulnerability wherein erased concepts can be reconstructed. This is achieved by identifying text tokens that closely align with the visual embeddings of the targeted concepts, thus enabling the regeneration of prohibited content [4, 72, 55, 59, 64]. This issue is illustrated in Fig. 1, demonstrating that even with concept erasure, T2I models can be manipulated through prompt modification to regenerate the restricted content. This underscores the imperative for a more robust concept erasure methodology that can withstand such reconstruction attempts, ensuring the integrity of content generation within T2I models.
@@ 3.2 Adversarial Training on Concept Erased Diffusion Models
To assess the efficacy of our method, we execute tests on a $\Phi_{\theta}$ model trained for removing the “nudity” concept via the ESD method. Utilizing 142 nudity-centric prompts from the I2P dataset [51], we systematically select timesteps $t^{*}$ at 100-step intervals within the $[1,1000]$ range and compute the Attack Success Rate (ASR), as formulated in Eq.(8), for each $t^{*}$. The results, depicted in Fig. 2, unveil a notable capability: the reconstruction of the erased concept is feasible even with the attack confined to a single timestep $t^{*}$. Interestingly, after $t^{*}=500$, we observe an increasing trend in the ASR. The modified images resulting from the attack visually convey the transition back to the erased concept and reflect the ASR patterns. For additional insights into various concepts and their corresponding ASR trends across different timesteps $t^{*}$, the supplementary material offers further information.
@@
Adversarial Training for T2I Concept Erasure. Motivated by the findings from our single-timestep adversarial attack experiments, we explore the potential of such attacks to enhance the robustness of concept erasure in T2I models, posing the question: Can adversarial attacks improve the resilience of concept erasure mechanisms?
@@ 4.1 Experimental Setting
Table 2: Attack Success Rate (ASR) against white (○)/black (●) box attacks. We conduct experiments on artistic style (Van Gogh), explicit concepts (nudity, violence, and illegal acts), and objects (church, golf ball, and parachute). “ESD/RACE-Concept” denotes the concept erased from the model. We also measure the quality of T2I models. We can observe that RACE reduces ASR by over 30% (-0.30 or below) in absolute value on Van Gogh, nudity, and church for UnlearnDiff [72], which is the previous SOTA attack method.
@@ 4.3 Disentanglement
Figure 4: RACE’s Disentanglement in Concept Erasure. This figure highlights RACE’s precision in erasing specific concepts, as shown in diagonal images, while preserving unrelated concepts, which is evident in off-diagonal images. For reference, baseline images generated by the original Stable Diffusion (SD) model are also presented.
Investigating RACE’s disentanglement performance, our study focuses on its capability to precisely erase intended concepts without impacting other elements. The evaluation spans both qualitative and quantitative measures. On the qualitative front, we configure separate RACE-enhanced models to specifically erase artistic imprints such as “Van Gogh”, “Thomas Kinkade”, and “Kilian Eng” from the SD. The illustrative outcomes, showcased in Fig. 4, underscore RACE’s erasure precision, ensuring that the excision of one artistic style doesn’t lead to the collateral removal of others. This meticulous erasure extends to discrete object concepts like “church”, “golf ball”, and “parachute”, with the generated images further affirming the method’s discernment.
@@ 4.4 Discussion
Table 4: Ablation Studies for Performance Improvement. Using the “nudity” concept, we evaluate the effectiveness of adding weight regularization and close concept keywords. ASR and quality metrics are measured to assess performance.
@@
Potential Strategy to Improve the Robustness-Quality Trade-off. Our findings underscore the ability of RACE to significantly bolster the SD model’s defense against prompt-based adversarial attacks across a variety of concepts. However, as illustrated in Tab.2, enhancing robustness appears to inversely impact image quality. To address this dichotomy, we test a refined version of the RACE loss function incorporating a regularization term:
@@ 5 Conclusion
In this work, we present RACE, a novel defense approach designed to protect the Text-to-Image Stable Diffusion models from prompt-based red teaming attacks. RACE effectively strengthens the model’s concept erasure capabilities while maintaining computational efficiency, offering a valuable enhancement to the current erasure framework and bolstering defenses against various adversarial techniques. We also observe the robustness-quality trade-off and discuss possible future directions to improve it. This initial contribution lays the groundwork for further exploration, underscoring the critical importance of developing sophisticated defenses in the rapidly evolving domain of generative AI.
Acknowledgements. We would like to express our gratitude to Sangmin Jung for assisting with the experiments. This work is partially supported by the National Science Foundation under Grant No. 2038666 and No. 2101052. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
Supplementary Material
Appendix 0.A Additional Related Works
0.A.1 Adversarial Training Approaches
@@
0.A.2 Additional Works in Concept Erase
@@
Appendix 0.B Additional Training Details
@@
Appendix 0.C Additional Analysis of Single-Timestep Adversarial Attack
In order to demonstrate the effectiveness of the proposed single-timestep adversarial attack, we perform additional experiments on $\Phi_{\theta}$ models, each trained to erase the “Van Gogh” and “church” concepts via the ESD method [10]. The results, as illustrated in Fig. 5, demonstrate that a single-timestep adversarial attack can successfully reconstruct previously erased concepts.
Appendix 0.D Extended Visual Results from RACE
In Figures Figs. 6, 7 and 8, we present a supplementary collection of images to further demonstrate the capability of our method in erasing targeted concepts. It is important to note that the set of images in Fig. 7 contains explicit content.
@@ Appendix 0.D Extended Visual Results from RACE
Appendix 0.E Determining PGD Hyperparameters
Within the main paper, we settle on $\epsilon=0.1$ and designate 10 steps for the adversarial attack. This section delves into the empirical analysis underpinning this selection, particularly focusing on the balance between ASR efficacy and hyperparameter tuning. As delineated in Tab. 5, an $\epsilon$ value of 0.2 marginally outperforms 0.1. However, as elaborated in the main manuscript, an elevated attack intensity with $\epsilon=0.2$ may compromise the equilibrium between adversarial robustness and the fidelity of generated images. Holding $\epsilon$ at 0.1, we observe a minimal variance between 10 and 20 attack steps, reinforcing our decision to configure PGD parameters at $\epsilon=0.1$ and 10 steps for an optimal trade-off.
@@ Appendix 0.E Determining PGD Hyperparameters
Appendix 0.F Challenges in Erasing Violence and Illegal Act
The data presented in Table 2 of the main manuscript indicate that “violence” and “illegal act” concepts exhibit less pronounced reductions in Attack Success Rate (ASR) against UnlearnDiff white box attack [72]. Specifically, “nudity” sees a 30% decrease in ASR, while “violence” and “illegal act” experience reductions of 11% and 5%, respectively. To align with the experimental framework of UnlearnDiff [72], we employed “violence” and “illegal act” as keywords for concept removal. It is postulated that these terms may not optimally represent the range of hazardous imagery producible by Stable Diffusion (SD) [46] through the prompts in the I2P dataset [51]. To examine this hypothesis, we calculate the CLIP-score [14] for explicit images generated by a baseline SD model, $SD(\tilde{y})$, where $\tilde{y}$ refers prompts from I2P and $c$ includes “nudity”, “violence”, and “illegal act” (i.e. CLIP-score($SD(\tilde{y})$,$c$)). As illustrated in LABEL:tab:keywㅌord-clipscore, a higher CLIP-score, indicating better keyword alignment with the dangerous images, correlates with a more substantial decrease in ASR. Conversely, keywords that poorly match the hazardous image content tend to result in less significant ASR reductions. This underscores the critical importance of selecting highly representative keywords for concept erasure in T2I models, as their alignment with the generated content significantly influences the effectiveness of concept erase.